# Learning responsibility allocations for multi-agent interactions: A differentiable optimization approach with control barrier functions

### v1

@@ Abstract
From autonomous driving to package delivery, ensuring safe yet efficient multi-agent interaction is challenging as the interaction dynamics are influenced by hard-to-model factors such as social norms and contextual cues. Understanding these influences can aid in the design and evaluation of socially-aware autonomous agents whose behaviors are aligned with human values. In this work, we seek to codify factors governing safe multi-agent interactions via the lens of responsibility, i.e., an agent’s willingness to deviate from their desired control to accommodate safe interaction with others. Specifically, we propose a data-driven modeling approach based on control barrier functions and differentiable optimization that efficiently learns agents’ responsibility allocation from data. We demonstrate on synthetic and real-world datasets that we can obtain an interpretable and quantitative understanding of how much agents adjust their behavior to ensure the safety of others given their current environment.
@@ II Related work
Recent works consider learning and instilling social awareness into autonomous planning algorithms to improve coordination between agents and, therefore, boost the overall safety and efficiency of the multi-agent system [6, 7, 8, 9]. The idea is to include another agent’s objective function into the autonomous agent’s planning objective. The weighting between the two objective terms, referred to as the social value orientation (SVO), determines the output behavior, ranging from prosocial, egotistic, altruistic, competitive, and masochistic. The SVO weightings can be learned from data [6] to understand how much an agent considers the welfare of others. Although these works do not frame their approach in terms of responsibility explicitly, the SVO weightings can potentially encode responsibility as they describe how much one agent considers the welfare of others. However, a few issues arise. First, the objectives can include other considerations beyond safety, making interpreting the SVO weightings challenging. Second, the SVO paradigm applies to only two agents, and it is unclear how it can extend to multiple agents. Third, it relies on knowledge of other agents’ reward functions, which is often hard to obtain, especially for multi-agent settings.
@@ III Problem formulation
Consider a system with $N$ agents where $x\in\mathcal{X}$ denotes the state describing all agents, $u=(u_{1},u_{2},...,u_{N})\in\mathcal{U}_{1}\times\mathcal{U}_{2}\times...\times\mathcal{U}_{N}$ denote the control inputs of each agent, and $\dot{x}=f(x,u)$ denotes the dynamics of the multi-agent system. Let $g(x,u)\geq 0$ denote a collision avoidance constraint shared amonst all agents. In real-world settings, humans take control actions to accomplish their desired task and avoid collision with one another. But how much does each agent contribute to ensuring the collision avoidance constraint is satisfied? In other words, how responsible is each agent in taking control actions that ensures that $g(x,u)\geq 0$ is satisfied. Suppose $\pi(\cdot)$ is a policy describing how humans behave around each other. How does $\pi$ encode each agent’s responsibility allocation in ensuring the collision avoidance constraint is satisfied?
While a human’s true behavior policy is extremely complex, we seek to develop an interpretable lens on $\pi$ to describe each agent’s responsibility allocation. Let $\boldsymbol{\gamma}$ denote the responsibility allocation for each agent. Then we seek to answer the following two questions: (i) What is an interpretable model describing the dependency of $\pi$ on $\boldsymbol{\gamma}$? (ii) Given such a model, how can $\boldsymbol{\gamma}$ be tractably inferred from human interaction data?
@@ IV-A Control barrier function (CBFs)
$\sup_{u\in\mathcal{U}}\>\nabla b(x)^{T}f(x,u)+\alpha(b(x))\geq 0,\quad\forall x\in\mathcal{X}.\vspace{-5mm}$ (1)
Intuitively, $b(\cdot)$ is a measure of safety, and a valid CBF ensures that there exists a feasible control preventing $b(x(t))$ from decreasing faster than a rate of $-\alpha(b(x))$ along system trajectories and that $\dot{b}(x(t))=0$ along the boundary of $\mathcal{C}$. It follows that if $b(\cdot)$ is a valid CBF for a set $\mathcal{C}$, then any locally Lipschitz controller $k:\mathcal{X}\rightarrow\mathcal{U}$ satisfying $\nabla b(x)^{T}f(x,u)+\alpha(b(x))\geq 0$ renders the set $\mathcal{C}$ forward invariant [17].
@@ IV-B CBF safety filters
A safety filter will take the desired control, $u^{\mathrm{des}}$, computed from an upstream planning/control module, and project it into the safe control set $\mathcal{U}^{\mathrm{safe}}(x;b,\alpha)$ defined in (2), i.e.
@@ IV-C Responsibility allocations via a CBF safety filter lens
In this work, we consider using CBF safety filters as models to describe multi-agent interactions and use them for system analysis and inference to gain insight into how agents cooperate to maintain safety. The insights gained into multi-agent interactions can be used for offline (crash) analysis, measuring social acceptability of a policy, or to adapt an agent’s behavior online.
Consider a multi-agent control affine dynamical system where $\mathbf{x}=(x_{1},...,x_{N})$ is the joint state describing $N$ agents, and each agent has control input $u_{i}$.
@@
Given a CBF $b(\cdot)$ for the multiagent system, Prob. 2 describes the direct application of the CBF safety filter problem (Prob. 1) onto the joint dynamics.
@@
$u^{\mathrm{exec}}_{1:N}=$ $\mathop{\mathrm{argmin}}_{u_{1},...,u_{N}}\quad\sum_{i=1}^{N}\|u_{i}-u^{\mathrm{des}}_{i}\|_{2}^{2}$
@@
$\>\quad\quad u_{1}\in\mathcal{U}_{1},...,u_{N}\in\mathcal{U}_{N}.$
Notice that Prob. 2 has equal weighting in penalizing each agent for deviating from their desired control input, implying no agent is more or less inclined to ensure the safety constraint is satisfied. We modify Prob. 2 by adding a coefficient $\gamma_{i}$ for each agent to influence the relative penalty in deviating away from the desired control. As such, we propose the following definition of responsibility allocation.
@@
Let there be $N$ agents interacting, and their joint dynamics are governed by (3). Let $b(\cdot)$ be a CBF for the joint system describing the collision set. Each agent $i\in\{1,2,\dots,N\}$ has a desired control input $u_{i}^{\mathrm{des}}$ that is projected into the safe control set via a projection mapping $\mathrm{proj}$. The projection mapping $\mathrm{proj}$ depends on the corresponding state $\mathbf{x}$, CBF $b$, class $\mathcal{K}_{\infty}$ function $\alpha$, and the responsibility allocation vector $\boldsymbol{\gamma}=[\gamma_{1},\gamma_{2},...,\gamma_{N}]$ with $\mathbf{1}^{T}\boldsymbol{\gamma}=1$ and $0\leq\gamma_{i}\leq 1$, which determines how much deviation from the desired control each agent is willing to make to satisfy (2). The projection mapping $\mathrm{proj}$ can be computed by solving the optimization problem described in Prob. 3 (differences from Prob. 2 are highlighted in blue).
@@
$\mathrm{s.t.}\quad\nabla b(\mathbf{x})^{T}\left[\tilde{f}(\mathbf{x})+\sum_{i=1}^{N}g_{i}(\mathbf{x})u_{i}\right]+\alpha(b(\mathbf{x}))\geq{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-\epsilon}$
$\qquad\>\>u_{1}\in\mathcal{U}_{1},...,u_{N}\in\mathcal{U}_{N}$
$\qquad\>\>{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\epsilon\geq 0}.$
@@
The slack variable $\epsilon>0$ reflects the discussion in Remark 2. The regularization term $\|u_{i}\|_{2}^{2}$ ensures a unique solution for cases in which Agent $i$ bears total responsibility for constraint satisfaction, i.e. $\gamma_{i}=0$.
@@
Let $\gamma$ denote the responsibility allocation of Agent 1. Fig. 2 shows the solution to Prob. 3 for various values of $\gamma$. We see that when $\gamma=0$, Agent 1 takes full responsibility by deviating from its desired control value to satisfy the CBF constraint (green region) while Agent 2 has no deivation. As $\gamma$ increases, Agent 2 begins to deviate more and Agent 1 less, until at $\gamma=1$, Agent 2 is fully responsible.
@@ V-A Bi-level optimization formulation
The goal is to find a value for $\boldsymbol{\gamma}$ such that the solution to Prob. 3 matches an interaction dataset as closely as possible. Suppose we have a dataset describing $N$-agent interactions $\mathcal{D}=\{(\mathbf{x}^{(k)},u_{1:N}^{(k)})\}_{k=1,...,K}$, and a desired policy $u^{\mathrm{des}}_{1:N}=\pi^{\mathrm{des}}_{1:N}(\mathbf{x})$ producing the agents’ desired control input. For example, each agent may wish to ignore the others and follow a nominal trajectory. Then, for a choice of CBF parameters $b$ and $\alpha$, we seek to find a responsibility allocation vector $\boldsymbol{\gamma}$ such that $u^{\mathrm{data}}_{1:N}\approx\mathrm{proj}(\pi_{1:N}^{\mathrm{des}}(\mathbf{x}^{\mathrm{data}});\mathbf{x}^{\mathrm{data}},b,\alpha,\boldsymbol{\gamma})$ for some choice of similarity loss metric $\Delta$ (e.g., L1, L2, Huber loss). Since the projection mapping $\mathrm{proj}$ contains an optimization problem (see Prob. 3), the resulting regression problem takes on a bi-level structure.
@@ V-B Leveraging differentiable optimization
1:$\widetilde{\boldsymbol{\gamma}}$ (unnormalized responsibility allocation)
2:$\delta$ (step size)
3:$\mathcal{D}=\{(\mathbf{x}^{(k)},u_{1:N}^{(k)})\}_{k=1,...,K}$ (dataset)
@@
5:$\widetilde{U}_{1:N}=\mathrm{batch}(\tilde{u}_{1:N}(\boldsymbol{\gamma},\mathbf{x}^{\mathrm{data}});\mathcal{D})$
6:$\ell(\boldsymbol{\gamma})=\sum_{(u_{1:N}^{\mathrm{data}},\tilde{u}_{1:N})\in\mathcal{D}\times\widetilde{U}_{1:N}}\Delta(u_{1:N}^{\mathrm{data}},\tilde{u}_{1:N})$
7:$\widetilde{\boldsymbol{\gamma}}\leftarrow\widetilde{\boldsymbol{\gamma}}-\delta\frac{\partial\ell}{\partial\boldsymbol{\gamma}}\frac{\partial\boldsymbol{\gamma}}{\partial\widetilde{\boldsymbol{\gamma}}}$
@@ V-C Context-dependent responsibility
Sec. V-B described the algorithm for finding the responsibility allocation parameter $\boldsymbol{\gamma}$ using gradient descent. That formulation assumed $\boldsymbol{\gamma}$ is fixed for each agent throughout the entire dataset. However, in general, the responsibility allocation is not agent-specific, but rather state- or context-dependent. Given that we are solving Prob. 4 using gradient descent, Alg. 1 can easily extend to the case where $\boldsymbol{\gamma}=\mathrm{softmax}(h_{\theta}(\mathbf{x},e))$ where $h_{\theta}:\mathcal{X}^{N}\times\mathcal{E}\rightarrow[0,1]^{N}$ is a differentiable function parameterized by $\theta$, such as a neural network, and depends on the state $\mathbf{x}$ and environment information $e$, such as map/road information. The $\mathrm{softmax}$ function ensures the $\boldsymbol{\gamma}$ constraints are satisfied. As such, we can perform gradient descent on the parameters $\theta$ instead.
@@
Figure 3: Learning $\gamma$ with synthetically generated data. After a few epochs, the estimated responsibility allocation value $\gamma$ converges to the ground truth value. The computation time scales linearly as the batch size increases.
@@ V-D Enforcing symmetry in responsibility allocation
For an $N$ agent system with joint state space $\mathcal{X}^{N}$, suppose some assignment ordering $(1,2,...,N)$. Let $\boldsymbol{\gamma}:\mathcal{X}^{N}\rightarrow[0,1]^{N}$ be the responsibility allocation function, and $\gamma_{i}(\mathbf{x})$ (with $\mathbf{x}=[x_{1},...,x_{N}]$) denote the responsibility allocation for agent $i$. Then, we say the system has symmetric responsibility allocation if the following properties hold:
@@
$\gamma_{i}(\mathbf{x})=\frac{\exp(\phi(\sigma_{i\rightarrow 1}(\mathbf{x})))}{\sum_{j=1}^{N}\exp(\phi(\sigma_{j\rightarrow 1}(\mathbf{x})))},\quad\phi(\mathbf{x})=\sum_{\sigma\in\mathcal{P}_{1}}\widetilde{\phi}(\sigma(\mathbf{x}))$ (7)
@@
Constructing a symmetric responsibility allocation function amounts to defining a single function $\widetilde{\phi}:\mathcal{X}^{N}\rightarrow\mathbb{R}$ and applying all possible $N-1$ permutations on it. As such, this construction can become prohibitively expensive as $N$ increases. In this work, we study cases with small $N$ values for which the computation is tractable, and defer investigating scalable approaches for large $N$ in future work.

### v2

@@ Abstract
From autonomous driving to package delivery, ensuring safe yet efficient multi-agent interaction is challenging as the interaction dynamics are influenced by hard-to-model factors such as social norms and contextual cues. Understanding these influences can aid in the design and evaluation of socially-aware autonomous agents whose behaviors are aligned with human values. In this work, we seek to codify factors governing safe multi-agent interactions via the lens of responsibility, i.e., an agent’s willingness to deviate from their desired control to accommodate safe interaction with others. Specifically, we propose a data-driven modeling approach based on control barrier functions and differentiable optimization that efficiently learns agents’ responsibility allocation from data. We demonstrate on synthetic and real-world datasets that we can obtain an interpretable and quantitative understanding of how much agents adjust their behavior to ensure the safety of others given their current environment. Project page: https://isremy.github.io/responsibility-website/
@@ II Related work
Recent works consider learning and instilling social awareness into autonomous planning algorithms to improve coordination between agents and thus boost the overall safety and efficiency of the multi-agent system [22, 25, 26, 24]. The idea is to include another agent’s objective function into the autonomous agent’s planning objective. The weighting between the two objective terms, referred to as the social value orientation (SVO), determines the output behavior, ranging from prosocial, egotistic, altruistic, competitive, and masochistic. The SVO weightings can be learned from data [22] to understand how much an agent considers the welfare of others. Although these works do not explicitly discuss responsibility, SVO weightings can potentially encode responsibility as they describe how much one agent considers the welfare of others. However, a few issues arise. First, the objectives can include other considerations beyond safety, making interpreting the SVO weightings challenging. Second, the SVO paradigm applies to only two agents, and it is unclear how it can extend to multiple agents. Third, it relies on knowledge of other agents’ reward functions, which is often hard to obtain, especially for multi-agent settings.
@@ III Problem formulation
Consider a system with $N$ agents where $x\in\mathcal{X}$ denotes the state describing all agents, $u=(u_{1},u_{2},...,u_{N})\in\mathcal{U}_{1}\times\mathcal{U}_{2}\times...\times\mathcal{U}_{N}$ denote the control inputs of each agent, and $\dot{x}=f(x,u)$ denotes the dynamics of the multi-agent system. Let $g(x,u)\geq 0$ denote a collision avoidance constraint shared amonst all agents. In real-world settings, humans take control actions to accomplish their desired task and avoid collision with one another. But how much does each agent contribute to ensuring the collision avoidance constraint is satisfied? In other words, how responsible is each agent in taking control actions that ensures that $g(x,u)\geq 0$ is satisfied. Suppose $\pi_{1:N}:\mathcal{X}\rightarrow\mathcal{U}_{1}\times\mathcal{U}_{2}\times...\times\mathcal{U}_{N}$ is an arbitrary policy describing how $N$ humans behave around each other. How does $\pi$ encode each agent’s responsibility allocation in satisfying collision avoidance constraints?
While a human’s true behavior policy is extremely complex, we seek to develop an interpretable lens on $\pi$ to describe each agent’s responsibility allocation. Let $\boldsymbol{\gamma}\in\mathbb{R}^{N}$ denote the responsibility allocation for each agent. Then we seek to answer the following two questions: (i) What is an interpretable model describing the dependency of $\pi$ on $\boldsymbol{\gamma}$? (ii) Given such a model, how can $\boldsymbol{\gamma}$ be tractably inferred from human interaction data?
@@ IV-A Control barrier function (CBFs)
$\sup_{u\in\mathcal{U}}\>\nabla b(x)^{T}f(x,u)+\alpha(b(x))\geq 0,\quad\forall x\in\mathcal{X}.\vskip-14.22636pt$ (1)
Intuitively, $b(\cdot)$ is a measure of safety, and a valid CBF ensures that there exists a feasible control preventing $b(x(t))$ from decreasing faster than a rate of $\alpha(b(x))$ along system trajectories and that $\dot{b}(x(t))=0$ along the boundary of $\mathcal{C}$. It follows that if $b(\cdot)$ is a valid CBF for a set $\mathcal{C}$, then any locally Lipschitz controller $k:\mathcal{X}\rightarrow\mathcal{U}$ satisfying $\nabla b(x)^{T}f(x,u)+\alpha(b(x))\geq 0$ renders the set $\mathcal{C}$ forward invariant [5].
@@ IV-B CBF safety filters
A safety filter will take the control from a desired policy, $u^{\mathrm{des}}=\pi^{\mathrm{des}}(x)$, and project it into the safe control set $\mathcal{U}^{\mathrm{safe}}(x;b,\alpha)$ defined in (2), i.e.
@@ IV-C Responsibility allocations via a CBF safety filter lens
In this work, we use CBF safety filters as models to describe multi-agent interactions, allowing us to infer how agents cooperate to maintain safety. The insights gained into multi-agent interactions can be used for offline (crash) analysis, measuring social acceptability of a policy, or to adapt an agent’s behavior online.
Consider a multi-agent control affine dynamical system where $\mathbf{x}=[x_{1},...,x_{N}]$ is the joint state describing $N$ agents, and each agent has control input $u_{i}$.
@@
Given a CBF $b(\cdot)$ for the multiagent system, Prob. 2 describes the direct application of the CBF safety filter problem (Prob. 1) onto the joint dynamics under $u^{\mathrm{des}}_{1:N}=\pi^{\mathrm{des}}_{1:N}(\mathbf{x})$.
@@
$u^{\mathrm{exec}}_{1:N}=$ $\mathop{\mathrm{argmin}}_{u_{1},...,u_{N}}\qquad\sum_{i=1}^{N}\|u_{i}-u^{\mathrm{des}}_{i}\|_{2}^{2}$
@@
$\>\qquad\qquad u_{1}\in\mathcal{U}_{1},...,u_{N}\in\mathcal{U}_{N}.$
Notice that Prob. 2 has equal weighting in penalizing each agent for deviating from their desired control input, implying no agent is more or less inclined to ensure the safety constraint is satisfied. We modify Prob. 2 by adding a coefficient $\boldsymbol{\gamma}_{i}$ for each agent to influence the relative penalty in deviating away from the desired control. As such, we propose the following definition of responsibility allocation.
@@
Let there be $N$ agents interacting, and their joint dynamics are governed by (3). Let $b(\cdot)$ be a CBF for the joint system describing the collision set. Each agent $i\in\{1,2,\dots,N\}$ has a desired control input $u_{i}^{\mathrm{des}}$ that is projected into the safe control set via a projection mapping $\mathrm{proj}$. The projection mapping $\mathrm{proj}$ depends on the corresponding state $\mathbf{x}$, CBF $b$, class $\mathcal{K}_{\infty}$ function $\alpha$, and the responsibility allocation vector $\boldsymbol{\gamma}=[\boldsymbol{\gamma}_{1},\boldsymbol{\gamma}_{2},...,\boldsymbol{\gamma}_{N}]$ with $\mathbf{1}^{T}\boldsymbol{\gamma}=1$ and $0\leq\boldsymbol{\gamma}_{i}\leq 1$, which determines how much deviation from the desired control each agent is willing to make to satisfy (2). The projection mapping $\mathrm{proj}$ can be computed by solving the optimization problem described in Prob. 3 (differences from Prob. 2 are highlighted in blue).
@@
$\mathrm{s.t.}\qquad\nabla b(\mathbf{x})^{T}\left[\tilde{f}(\mathbf{x})+\sum_{i=1}^{N}g_{i}(\mathbf{x})u_{i}\right]+\alpha(b(\mathbf{x}))\geq{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-\epsilon}$
$\hskip 18.49988pt\>\>u_{1}\in\mathcal{U}_{1},...,u_{N}\in\mathcal{U}_{N}$
$\hskip 18.49988pt\>\>{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\epsilon\geq 0}.$
@@
The slack variable $\epsilon>0$ reflects the discussion in Remark 2. The regularization term $\|u_{i}\|_{2}^{2}$ ensures a unique solution for cases in which Agent $i$ bears total responsibility for constraint satisfaction, i.e. $\boldsymbol{\gamma}_{i}=0$. $\beta_{1,2}$ are scalar hyperparameters for regularization and $\epsilon$ weighting.
@@
Let $\gamma\in[0,1]$ denote the responsibility allocation of Agent 1. Fig. 2 shows the solution to Prob. 3 for various values of $\gamma$. We see that when $\gamma=0$, Agent 1 takes full responsibility by deviating from its desired control value to satisfy the CBF constraint (green region) while Agent 2 has no deviation. As $\gamma$ increases, Agent 2 begins to deviate more and Agent 1 less, until at $\gamma=1$, Agent 2 is fully responsible.
@@ V-A Bi-level optimization formulation
The goal is to find a value for $\boldsymbol{\gamma}$ such that the solution to Prob. 3 matches an interaction dataset as closely as possible. Suppose we have a dataset describing $N$-agent interactions $\mathcal{D}=\{(\mathbf{x}^{(k)},u_{1:N}^{(k)})\}_{k=1,...,K}$, and a desired policy $\pi^{\mathrm{des}}_{1:N}(\mathbf{x})$ producing the agents’ desired control inputs $u^{\mathrm{des}}_{1:N}$. For example, each agent may wish to ignore the others and follow a nominal trajectory. Then, for a choice of CBF parameters $b$ and $\alpha$, we want a responsibility allocation vector $\boldsymbol{\gamma}$ such that $u^{\mathrm{data}}_{1:N}\approx\mathrm{proj}(\pi_{1:N}^{\mathrm{des}}(\mathbf{x}^{\mathrm{data}});\mathbf{x}^{\mathrm{data}},b,\alpha,\boldsymbol{\gamma})$ for some distance metric $\Delta$ (e.g., L1, L2, Huber loss). Since the projection mapping $\mathrm{proj}$ is an optimization problem (see Prob. 3), the resulting regression problem has a bi-level structure.
@@ V-B Leveraging differentiable optimization
[…]
@@ V-C Context-dependent responsibility
Sec. V-B described the algorithm for finding the responsibility allocation parameter $\boldsymbol{\gamma}$ using gradient descent. That formulation assumed $\boldsymbol{\gamma}$ is a fixed vector describing agent responsibilities for an entire dataset. However, in general, the responsibility allocation is not agent-specific, but rather state- or context-dependent (i.e., $\boldsymbol{\gamma}$ is a function). Given that we are solving Prob. 4 using gradient descent, Alg. 1 can easily extend to the case where $\boldsymbol{\gamma}=\mathrm{softmax}(h_{\theta}(\mathbf{x},e))$ where $h_{\theta}:\mathcal{X}^{N}\times\mathcal{E}\rightarrow[0,1]^{N}$ is a differentiable function parameterized by $\theta$, such as a neural network, and depends on the state $\mathbf{x}$ and environment information $e$, such as map/road information. The $\mathrm{softmax}$ function ensures the $\boldsymbol{\gamma}$ constraints are satisfied. As such, we can perform gradient descent on the parameters $\theta$ instead.
@@
Figure 3: We generate ground truth $\gamma$ values to create synthetic $u_{1:N}$ via Prob. 3, then learn the original $\gamma$ values via Prob. 4. After a few epochs, the estimated responsibility allocation value $\gamma$ converges to the ground truth. Computation time scales linearly with batch size.
@@ V-D Enforcing symmetry in responsibility allocation
For an $N$ agent system with joint state space $\mathcal{X}^{N}$, suppose some assignment ordering (1,2,,N)(1,2,...,N). Let $\boldsymbol{\gamma}:\mathcal{X}^{N}\rightarrow[0,1]^{N}$ be the responsibility allocation function, and $\gamma_{i}(\mathbf{x})$ (with $\mathbf{x}=[x_{1},...,x_{N}]$) denote the responsibility allocation for agent $i$. Then, we say the system has symmetric responsibility allocation if the following properties hold:
@@
$\gamma_{i}(\mathbf{x})=\frac{\exp(\phi(\sigma_{i\rightarrow 1}(\mathbf{x})))}{\sum_{j=1}^{N}\exp(\phi(\sigma_{j\rightarrow 1}(\mathbf{x})))},\qquad\phi(\mathbf{x})=\sum_{\sigma\in\mathcal{P}_{1}}\widetilde{\phi}(\sigma(\mathbf{x}))$ (7)
@@
Constructing a symmetric responsibility allocation function amounts to defining a single function $\widetilde{\phi}:\mathcal{X}^{N}\rightarrow\mathbb{R}$ and applying all possible N1N-1 permutations on it. As such, this construction can become prohibitively expensive as $N$ increases. In this work, we study cases with small $N$ values for which the computation is tractable, and defer investigating scalable approaches for large $N$ in future work.